OpenStudy (anonymous):

Urgent, help please! I buy some small bags of chips and some big bag of chips. The small bag weighs 150g and he big bags weigh 750 g. The total number of bags of chips is 16 and their total wight is 6 kg. How many small bags and how my big bags did I buy? PLEASE HELP!

OpenStudy (anonymous):

solve this system x+5y=40 x+y=16

OpenStudy (anonymous):

Where did you get 5 and 40 from?

OpenStudy (anonymous):

150x+750y=6000

OpenStudy (anonymous):

I dont get it.. can you do the whole math? please? i dont have much time! please please pretty please!

OpenStudy (anonymous):

x bag of 150g packet weigh 150x g y bag of 750g packet weigh 750y g total weight=6000g 150x+750y=6000

OpenStudy (anonymous):

then have to find the x and y?

OpenStudy (anonymous):

Y